OpenStudy (anonymous):

Megan bought 7 charms for $31.50. Each charm costs the same amount of money. Whats an inequality that can be used to find the maximum amount of charms Megan can buy with $75? And what's the maximum amount of charms she can buy too?

OpenStudy (anonymous):

CAN SOMEONE HELP ME

OpenStudy (anonymous):

With what? I asked the question lol

OpenStudy (anonymous):

\[7charms=31.50\] \[31.50\div7=4.5\] So, each charm is $4.50, and let x represent the amount of charms; \[4.5x \le75\] This shows that $4.50 per charm has to be less than or equal to $75. Now all you have to do to solve it is divide both sides by 4.5; \[\frac{ 4.5x }{ 4.5 }\le \frac{ 75 }{ 4.5 }\] So, \[x \le16.66...\] But you can't have .6666 of a charm, so the amount of charms she can buy with $75 is 16 charms.

OpenStudy (anonymous):

HEY I4NI CAN YOU HELP ME ??

OpenStudy (anonymous):

I can try, make another question topic and I will see what I can do.

OpenStudy (anonymous):

You (or your parents) earn $53,698.00/yr and you have deductions of FICA (7.65%), federal tax withholding (11.5%), and state tax withholding (9%). How much is the allowed housing expense for the year? (1 point) $ /year 4. You (or your parents) plan to pay $1,275.00/month for a mortgage. How much is the minimum realized income per month to the nearest penny

OpenStudy (anonymous):

You are incredible I4NI Thanks for the help, I think i get it now=)

OpenStudy (anonymous):

No problem (: Trapp, make another topic for your question, as this one was for the OP's question.

OpenStudy (anonymous):

I DID , I JUST HAVE TO WAIT 1 MINUTE TO BUMP IT UP !

OpenStudy (anonymous):

Well than wait one minute=)